• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ransome

    Log entry time 23:44:36 on July29,2006

    Entry number 176118

    keyword=Shift summary swing July 29

    Begin shift with BCM calibration underway
    Finished 16:20
    Move 15 cm LD2 target in, begin ed data taking.
    1833 ed - mainly at 7 uA, beam recentered (was off .4 mm),
    increased to 10 uA - 800K
    1834 ed - event rate ~1 kHz, dead time is 15% - 1 million
    1835 ed - 1 million
    1836 ed - 1 miillion
    1837 ed - 1 miilion
    1838 ed - 1 million
    1839 ed - 1 million
    1840 ed - 1 million
    1841 ed - 1 million
    1842 ed - 1 million - total 9.8 million finished 19:50
    switch to 4% radiator
    1843 gamma d - started - 2 million event
    Half wave plate IN now
    1844 gamma d - 2 million
    1845 gamma d - 2 million
    1846 gamma d - 2 million
    1847 gamma d - 2 million
    1848 gamma d - 2 million
    1849 gamma d - 2 million
    1850 gamma d - 2 million
    1851 gamma d - 2 million
    1852 gamma d - 2 million
    1853 gamma d - 2 million
    1854 gamma d - 2 million
    1855 gamma d - 2 million
    A good shift